# definition

### v2

@@ E.3 Proof of Lemma 12 (Sufficient coverage)
$\overline{C}^{*}([N(m({\bm{X}}_{t})+1/N)])$ $=\overline{C}^{*}([Nm({\bm{X}}_{t})+1])-\overline{C}^{*}([N])$
@@
Appendix F Proofs of Lemma 4 and Lemma 5
In this section, we provide two final lemmas, Lemma 4 and Lemma 5, which together imply Theorem 1. We prove these two lemmas in Sections F.1 and F.2, respectively.
F.1 Proof of Lemma 4
@@
$=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\Big{(}R^{\textup{rel}}(N,\bm{S}_{0})-\frac{1}{N}\sum_{i\in[N]}\mathbb{E}[r_{i}(S_{i,t},A_{i,t})]\Big{)}$
@@
$=\frac{1}{N}\sum_{i\in[N]}\left\langle r_{i}^{*},\mu^{*}_{i}-\mathbb{E}[X_{i,t}]\right\rangle+\frac{2r_{\max}}{N}\sum_{i\in[N]}\mathbb{E}\left[\mathbf{1}\left\{A_{i,t}\neq\widehat{A}_{i,t}\right\}\right]$
$\leq\frac{1}{N}\sum_{i\in[N]}\left\langle r_{i}^{*},\mu^{*}_{i}-\mathbb{E}[X_{i,t}]\right\rangle+2r_{\max}\mathbb{E}\left[1-\frac{N_{t}^{*}}{N}\right]$ (75)
$\leq\frac{1}{N}\sum_{i\in[N]}\left\langle r_{i}^{*},\mu^{*}_{i}-\mathbb{E}[X_{i,t}]\right\rangle+2r_{\max}\mathbb{E}\left[1-m({\bm{X}}_{t})\right]+\frac{K_{\text{conf}}}{\sqrt{N}}$ (76)
$\leq\frac{1}{N}\mathbb{E}\left[h_{\textnormal{ID}}({\bm{X}}_{t},1)\right]+2r_{\max}\mathbb{E}\left[1-m({\bm{X}}_{t})\right]+\frac{K_{\text{conf}}}{\sqrt{N}}$ (77)
@@
$h_{\textnormal{ID}}({\bm{X}}_{t},1)$ $\leq h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))+L_{h}N(1-m({\bm{X}}_{t}))=V({\bm{X}}_{t}),$ (78)
$1-m({\bm{X}}_{t})$ $\leq\frac{1}{L_{h}N}V({\bm{X}}_{t})$ (79)
@@
$R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})$ $\leq\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\Big{(}\frac{1}{N}\mathbb{E}\left[h_{\textnormal{ID}}({\bm{X}}_{t},1)\right]+2r_{\max}\mathbb{E}\left[1-m({\bm{X}}_{t})\right]+\frac{K_{\text{conf}}}{\sqrt{N}}\Big{)}$
$\leq\frac{2r_{\max}+L_{h}}{L_{h}N}\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[V({\bm{X}}_{t})\right]+\frac{K_{\text{conf}}}{\sqrt{N}}\,.$
F.2 Proof of Lemma 5
@@
$\mathbb{E}[h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))\;|\;{\bm{X}}_{t}]\leq\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))+\left(2C_{h}+2\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}\right)\sqrt{N}\,.$ (81)
@@
$+\left(2C_{h}+2\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}+2L_{h}K_{\text{mono}}\right)\sqrt{N}.$
@@
$K_{V}$ $=2C_{h}+2\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}$
$+2L_{h}K_{\text{mono}}+\frac{\rho_{V}L_{h}K_{\text{cov}}}{\sqrt{N}}.$
@@
which completes the proof.
Finally, we give a more explicit form of the constant $C_{\textnormal{ID}}$ in Theorem 1 based on the proof above. Combining the optimality gap bound in Lemma 4 with the inequality in (85), we get
$R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})\leq\frac{(2r_{\max}+L_{h})K_{V}}{L_{h}\rho_{V}\sqrt{N}}+\frac{K_{\text{conf}}}{\sqrt{N}}.$
We have thus proved $R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})\leq C_{\textnormal{ID}}/\sqrt{N}$, where $C_{\textnormal{ID}}$ is independent of $N$ and is given by
$C_{\textnormal{ID}}$ $=\frac{\left(2r_{\max}+L_{h}\right)\left(1/L_{h}+1/\eta_{c}\right)}{1-\gamma}\left(2C_{h}+2\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}+2L_{h}K_{\text{mono}}\right)$
$+2(r_{\max}+L_{h})K_{\text{cov}}+K_{\text{conf}}\,.$

### v3

@@ F.3 Proof of Lemma 10 (Sufficient coverage)
$\overline{C}^{*}([Nm({\bm{X}}_{t})+1])$ $\geq\overline{C}^{*}([Nm({\bm{X}}_{t})+1])-\overline{C}^{*}([N])$
@@
Appendix G Proofs of Lemma 4 and Lemma 5
In this section, we provide two final lemmas, Lemma 4 and Lemma 5, which together imply Theorem 1. We prove these two lemmas in Sections G.1 and G.2, respectively.
G.1 Proof of Lemma 4
@@
$=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\Big{(}R^{\textup{rel}}(N,\bm{S}_{0})-\frac{1}{N}\sum_{i\in[N]}\mathbb{E}[r_{i}(S_{i,t},A_{i,t})]\Big{)}.$
@@
$=\frac{1}{N}\sum_{i\in[N]}\left\langle\mu^{*}_{i}-\mathbb{E}[X_{i,t}],r_{i}^{*}\right\rangle+\frac{2r_{\max}}{N}\sum_{i\in[N]}\mathbb{E}\left[\mathbf{1}\left\{A_{i,t}\neq\widehat{A}_{i,t}\right\}\right]$
$\leq\frac{1}{N}\sum_{i\in[N]}\left\langle\mu^{*}_{i}-\mathbb{E}[X_{i,t}],r_{i}^{*}\right\rangle+2r_{\max}\mathbb{E}\left[1-\frac{N_{t}^{*}}{N}\right]$ (72)
$\leq\frac{1}{N}\sum_{i\in[N]}\left\langle\mu^{*}_{i}-\mathbb{E}[X_{i,t}],r_{i}^{*}\right\rangle+2r_{\max}\mathbb{E}\left[1-m({\bm{X}}_{t})\right]+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}$ (73)
$\leq\frac{1}{N}\mathbb{E}\left[h_{\textnormal{ID}}({\bm{X}}_{t},1)\right]+2r_{\max}\mathbb{E}\left[1-m({\bm{X}}_{t})\right]+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}},$ (74)
@@
$h_{\textnormal{ID}}({\bm{X}}_{t},1)$ $\leq h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))+L_{h}N(1-m({\bm{X}}_{t}))=V({\bm{X}}_{t})$ (75)
$1-m({\bm{X}}_{t})$ $\leq\frac{1}{L_{h}N}V({\bm{X}}_{t}),$ (76)
@@
$R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})$ $\leq\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\Big{(}\frac{1}{N}\mathbb{E}\left[h_{\textnormal{ID}}({\bm{X}}_{t},1)\right]+2r_{\max}\mathbb{E}\left[1-m({\bm{X}}_{t})\right]+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}\Big{)}$
$\leq\frac{2r_{\max}+L_{h}}{L_{h}N}\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[V({\bm{X}}_{t})\right]+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}\,.\qed$
G.2 Proof of Lemma 5
@@
$\mathbb{E}[h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))\;|\;{\bm{X}}_{t}]\leq\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))+\big{(}2+K_{\text{conf}}\big{)}C_{h}\sqrt{N}\,.$ (78)
@@
$+\left(2C_{h}+C_{h}K_{\text{conf}}+L_{h}K_{\text{mono}}\right)\sqrt{N}.$
@@
$K_{V}=2C_{h}+C_{h}K_{\text{conf}}+2L_{h}K_{\text{mono}}+\frac{\rho_{V}L_{h}K_{\text{cov}}}{\sqrt{N}}.$
@@
which completes the proof. ∎
Finally, we give a more explicit form of the constant $C_{\textnormal{ID}}$ in Theorem 1:
  • If $c_{\max}<\alpha_{\min}$, the budget constraints are vacuous and all arms will follow the ideal actions under the ID policy. Consequently, $R^{*}(N,\bm{S}_{0})=R(\pi,\bm{S}_{0})$ and we can simply take $C_{\textnormal{ID}}=0$.
  • If $c_{\max}\geq\alpha_{\min}$, we combine the optimality gap bound in Lemma 4 with (82) to get
    $R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})\leq\frac{(2r_{\max}+L_{h})K_{V}}{L_{h}\rho_{V}\sqrt{N}}+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}.$
    We therefore obtain $R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})\leq C_{\textnormal{ID}}/\sqrt{N}$, where $C_{\textnormal{ID}}$ is given by
    $C_{\textnormal{ID}}$ $=\frac{\left(2r_{\max}+L_{h}\right)\left(1/L_{h}+1/\eta_{c}\right)}{1-\gamma}\left(2C_{h}+C_{h}K_{\text{conf}}+2L_{h}K_{\text{mono}}\right)$
    $+(2r_{\max}+L_{h})K_{\text{cov}}+2r_{\max}K_{\text{conf}}$
    $=O\left(\frac{K^{5}\max\{r_{\max},c_{\max}\}^{7}\tau^{4}}{\alpha_{\min}^{6}}\right).$
    In the last step above, we substituted in all intermediate constants from the proofs; one can verify that each of these constants depends on the problem parameters as follows:
    $L_{h}$ $=O\big{(}\max\{r_{\max},c_{\max}\}\tau\big{)}$
    $C_{h}$ $=O\big{(}K\max\{r_{\max},c_{\max}\}\tau\big{)}$
    $\frac{1}{1-\gamma}$ $=O(\tau)$
    $\frac{1}{\eta_{c}}$ $=O\left(\frac{Kc_{\max}}{\alpha_{\min}^{2}}\right)$
    $K_{\text{conf}}$ $=O\left(\frac{K^{2}c_{\max}^{2}}{\alpha_{\min}^{2}}\right)$
    $K_{\text{mono}}$ $=O\left(\frac{K^{4}c_{\max}^{3}\max\{r_{\max},c_{\max}\}}{\alpha_{\min}^{4}}\tau\right)$
    $K_{\text{cov}}$ $=O\left(\frac{Kc_{\max}\max\{r_{\max},c_{\max}\}}{\alpha_{\min}^{2}}\tau\right).$